OpenStudy (anonymous):

HELP! R=KL/d2 (d squared)

OpenStudy (campbell_st):

what do you need to do..?

OpenStudy (anonymous):

oh solve for d

OpenStudy (campbell_st):

ok... multply both sides by d^2 which will get d out of the denominator

OpenStudy (anonymous):

and i would get R(d^2)=KL?

OpenStudy (campbell_st):

you now divide both sides by R

OpenStudy (campbell_st):

last step will be taking the square root of everything on the right hand side...

OpenStudy (anonymous):

d= the square root of KL/R ?

OpenStudy (campbell_st):

yes... now I'm not sure how musch you have done on taking square roots... but there is actually 2 answers... normally written \[d = \pm \sqrt{\frac{KL}{r}}\]

OpenStudy (campbell_st):

oops should be R

OpenStudy (anonymous):

ohhh I see . thanks . i have another one .

OpenStudy (anonymous):

\[\frac{ V _{p} }{ V _{s} }=\frac{ N _{p} }{ N _{s} } solving for N _{s}\]

OpenStudy (campbell_st):

cross multiply then divide both sides by Vp

OpenStudy (anonymous):

 ?